A stock has a beta of 1.5 and an expected return of 13.3 percent. If the risk-free rate is 4.1 percent, what is the market risk premium?(Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Market risk premium %
Answer #1

E(r) = Rf + (Rm-Rf)*Beta

13.3 = 4.1 + (Rm - 4.1)1.5

13.3 = 4.1 + 1.5Rm - 6.15

15.35 = 1.5Rm

Rm = 10.23%

Market risk premium = 10.23%